# A Centralized Planning and Distributed Execution Method for Shape Filling with Homogeneous Mobile Robots

### v1

@@ B.7 Conventions and Notations
  • robot $u$ belongs to ribbon $R$ iff its center $p_{n^{-}}(u)$ is a vertex of $R$, denoted as $u\in R$. The gradient value of robot $u$ is the gradient value of the ribbon $R$;
  • robot $u$ is in front of robot $v$ iff $u,v$ belongs to the same ribbon and lattice point $p_{n^{-}}(u)$ is in front of lattice point $p_{n^{-}}(v)$ with respect to the ribbon.
@@
  • a ribbon $R$ is filled iff $\forall x\in V(R),\exists\text{non-moving robot}u\in U$ such that $x=p_{n^{-}}(u)$;
@@ Appendix C Correctness of the Add-subtract Algorithm
In appendix C, we formalize the add-subtract algorithm and provide a rigorous proof of correctness. Suppose we have obtained a proper ribbonization of $S$ and suppose there are $N$ lattice points in $S$.
Firstly, we define the activation sequence $s[k]$ and the assembly sequence $t[k]$ in section C.1. In order to prove that the robots fill $S$ after the additive stage, we investigate the mobility and localizability of each active robot of the additive stage. In section C.2, we show the mobility of each moving robot $u_{n}$ by investigating the boundary of the MRS. In particular, we prove that the boundary of the MRS maintains a certain smoothness that allows an unobstructed path along the perimeter of the MRS from start position $s[k]$ to target position $t[k]$ for every active robot $u_{k}$ with $k\leq N$. To ensure the algorithm can be implemented distributively, we show the localizability of each moving robot $u_{n}$ in section C.3. In particular, we show that the active robot is able to localize itself at the starting and stopping positions. Since the path is unobstructed and the robot is localizable, we conclude that robot $u_{k}$ stops at $t[k]$ (theorem 19).
Secondly, we investigate the properties of the MRS after the additive stage in section C.4. In particular, we define the "merging point" for every ribbon to denote the lattice point on the perimeter at which a robot moves out of $S$ first visits and the properties of the "merging point" are studied.
Thirdly, we formalize the subtractive stage of the algorithm by defining the reactivation sequence $s_{r}[k]$ and reassembly sequence $t_{r}[k]$ (see section C.5). In order to show that the extra robots of each ribbon are correctly removed and the remaining robots fill up $S\setminus D$, we need to prove the mobility and localizability of each active robot for the subtractive stage. In section C.6, we investigate the path between lattice point $s_{r}[k]$ and lattice point $t_{r}[k]$, and show that the path for robot $u_{N+k}$ is unobstructed by utilizing the results from the additive stage. To ensure the algorithm can be implemented distributively, we prove the localizability of each moving robot in section C.7. Since the path is unobstructed and the robot is localizable, we conclude that robot $u_{N+k}$ moves from $s_{r}[k]$ and stops at $t_{r}[k]$ (theorem 34).
Finally, in section C.8, we conclude that the movement sequences defined in the add-subtract algorithm satisfies the problem statement.
@@ C.1 Activation Sequence and Assembly Sequence
During the additive stage, in each epoch, one of the idle robots become active and moves along the perimeter of the collection of the robots, the moving robot enters the user-defined shape and stops at one of the lattice point inside the shape. In such a way, the MRS shrinks in the idle half-plane $H_{2}$ and expends inside $S$ in the formation half-plane $H_{1}$. We use two sequences of lattice points to represent which idle robot is activated and where it shall stop:
@@ C.2 Path for the Additive Stage
In the similar way, we define the edge-following points and the corresponding edge-following set $\bm{EFS_{n^{+}}}$ for sub-epoch $T=n^{+}$.
@@
We shall prove by induction. Let $P(n)$ be "at sub-epoch $T=n^{-}$, the set of boundary robots $B_{n^{-}}$ is proper". Let $Q(n)$ be "at sub-epoch $T=n^{+}$, the set of boundary robots $B_{n^{+}}$ is proper".
@@
        Case 3: $v$ is an idle robot. Since any lattice point in $H_{2}$ is at least $2\sqrt{3}d$ from any lattice point in $H_{1}$, any robot occupying a lattice point in $H_{1}$ and an idle robot in $H_{2}$ cannot be adjacent to a same lattice point.
@@
        Case 2 & Case 3: $v$ is a seed robot or an idle robot. By the similar arguments as in 1.Case 2 and 1.Case 3, we obtain contradictions.
@@
Since an edge-following path consists of arcs between adjacent edge-following points, we could define a robots’ contribution to the path as follows:
@@ C.3 Localization for the Additive Stage
We shall do induction on epoch $T$. Let $P(n)$ be "$u_{n}$ is localizable if it neighbors a seed robot or a robot that belongs to a filled ribbon."
@@
Theorem 19.
At epoch $1\leq T\leq N$, robot $s_{n}$ executing the edge-following movement stops at $t[n]$.
Proof.
By lemma 15 and lemma 18, the theorem follows. ∎
@@ C.4 Properties of the MRS At the End of the Additive Stage
    If $x$ has the type $A$ neighborhood, denote the lattice point in the neighborhood with a hop count $k-1$ as $y$. Two lattice points in the neighborhood of $x$ are adjacent to both $x$ and $y$, and we denote the lattice point on the counter-clockwise side of $x$ with respect to $y$ as $a$, and the lattice point on the clockwise side of $x$ as $b$. Since $R$ has a minimum length of two, $a$ is a vertex of $R$ and $a$ is behind $x$. As a result, $b$ is not in $S$, and $b$ is a merging point of $R$. Now we show that $b$ is the only merging point of $R$.
@@
    Inductive step: suppose $Q(n-1)$ is true, we want to show $P(n)$ is true; and suppose $P(n)$ is true, we want to show that $Q(n)$ is true. Since $s[n]$ is in the idle half-plane $H_{2}$, the non-moving robots in $H_{1}$ and the affiliated path remains unchanged after $u_{n}$ starts to move. As a result, $Q(n-1)\rightarrow P(n)$. We are left to show $P(n)\rightarrow Q(n)$. In other words, we want to show all the points contributed by robots of $R$ are behind $m$ after $u_{n}$ stops. We shall discuss the following three cases:
@@
First note that $m$ is on the culminating edge-following path $\widehat{EFP}$ since $m$ is adjacent to a non-moving robot in $S$ at sub-epoch $T=N^{+}$. As a result, $m$ is on $\widehat{AP}$. By proposition 21, the proposition follows.
@@
First notice that by the definition of merging point, $n$ is contributed by a robot of $R_{i}$. By proposition 22, $m$ is in front of all the points on the affiliated path that are contributed by robots from $R_{i}$. The corollary follows. ∎
@@ C.5 Reactivation Sequence and Reassembly Sequence
During the subtractive stage, extra robots of each layer move outside $S$ while others re-arrange themselves along the layer to fill up the lattice points in $S\setminus D$. The robots moving out of $S$ move along the perimeter of the MRS and join the idle robots in the idle half-plane $H_{2}$. We call such robots as "recycled robots". The remaining robots of each layer move within the layer and redistribute to occupy lattice points in $S\setminus D$, and we call such robots "rearranged robots".
@@
The reactivation sequence $s_{r}[k]$ is defined recursively such that each element $s_{r}[k]$ is the largest lattice point in $S$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${s_{r}}_{[1:k-1]}$.
@@
A recycle robot $u$ will exit the shape $S$ and stop in the idle half-plane $H_{2}$; and a rearranged robot $u$ will reposition itself inside $S\setminus D$.
@@
    The rearrangement sequence is a sequence of lattice points in $S\setminus D$, defined recursively such that the $k$th element is the largest lattice point in $S\setminus D$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not yet been included in the rearrangement sequence.
@@
The reassembly sequence $t_{r}$ is defined such that:
  1. 1.
    if $ID(s_{r}[1])<c(R(s_{r}[1]))$, $t_{r}[1]$ is the first lattice point in the recycle sequence;
  2. 2.
    if $ID(s_{r}[1])\geq c(R(s_{r}[1]))$, $t_{r}[1]$ is the first lattice point in the rearrangement sequence.
and for $k>1$,
  1. 1.
    if $ID(s_{r}[k])<c(R(s_{r}[k]))$, $t_{r}[k]$ is the next lattice point in the recycle sequence;
  2. 2.
    if $ID(s_{r}[k])\geq c(R(s_{r}[k]))$, $t_{r}[k]$ is the next lattice point in the rearrangement sequence.
Loosely speaking, the robot at the front of the outermost ribbon becomes active first. If the robot is a recycled robot, it moves out of the shape $S$ and stops in $H_{2}$ layer by layer; otherwise the robot (rearranged robot) moves to the front of the ribbon and stops at the next unoccupied vertex. This property is captured by the following proposition:
@@
Let sequence $I[k]$ be an index sequence of selects the first vertex of each ribbon in the reactivation sequence $s_{r}$. Notice that $I[k]$ follows the complete tree order in the sense that each element appears earlier in the sequence $I[k]$ is smaller with respected to the complete tree order. Let $Q$ denotes the set containing all boundary points of $S$.
@@ C.6.1 Motion of Recycled Robots
For a recycle robot $u_{n}$ with $n=N+k$, suppose $s_{r}[k]$ is a vertex of ribbon $R$ and $R$ has a parent ribbon $R^{[p]}$. The active robot first rotates around the parent ribbon $R^{[p]}$ while keeping distance $d$ from the nearest robot on $R^{[p]}$ until it reaches the first vertex of $R$. From the first vertex $x$, the robot continuous its rotation and reaches the merging point $m(R)$ of the ribbon $R$. Finally, $u_{n}$ moves clockwise along the perimeter of the MRS and stops in $H_{2}$.
@@
Notice that all lattice points in $R$ before $s_{r}[k]$ are unoccupied since all robots occupied such lattice points have moved out of $S$ and stopped in $H_{2}$ by proposition 24. It is left to show that the curve is unobstructed between each adjacent lattice point on the path. Suppose two adjacent lattice point $x,y$ are vertices of ribbon $R$, the curve between $x$ and $y$ with distance $d$ from robot $v$ (belonging to the parent ribbon of $R$) is shown below.
@@
The extended edge-following set, denoted as $\overline{EFS_{n^{-}}}$, is the set that contains all the extended edge-following points. The same definition applies to the extended edge-following points at $T=n^{+}$ and corresponding extended edge-following set $\overline{EFS_{n^{+}}}$.
@@
We shall prove by induction. Let $P(n)$ be "the extended set of boundary robots $\overline{B_{n}^{-}}$ is proper". Let $Q(n)$ be "the extended set of boundary robots $\overline{B_{n}^{+}}$ is proper".
First, notice the culminating set of boundary robots $\widehat{B}$ is proper. Since the extended set of boundary robots agrees with the culminating set of boundary robots in $H_{1}$, the base case $P(N+1)$ is true.
@@
Recall $p_{3}$ is the lattice point on the right of the origin $O$. Observe that the exterior edge-following path goes through lattice point $p_{3}$. Since virtual robots (if exist) are in half-plane $H_{1}$, thus we only need to show the path segment of the exterior edge-following path from $m(R)$ to $p_{3}$ satisfies the proposition. We denote the path segment from $m(R)$ to $p_{3}$ as $P$. Notice that $P$ is a path segment of the culminating affiliated path $\widehat{AP}$.
By proposition 25, no point on $P$ is contributed by any virtual robot at the time when the first robot of ribbon $R$ starts to move. Let $T=l$ denote the epoch when the first robot of ribbon $R$ start to move. By the reactivation sequence, all active robots at epoch $l<T<N+k$ are from ribbon $R$. As a result, a point on $P$ is either contributed by a boundary robot or a virtual robot occupying some vertex of $R$. By proposition 22, $m(R)$ is in front of all points of $\widehat{AP}$ that are contributed by robots/virtual robots of $R$. As a result, no point on $P$ (except for the merging point itself) is contributed by virtual robot. The proposition follows.
@@
For recycle robot $u_{n}$ with $n=N+k$, its exterior path is well-defined and unobstructed.
@@ C.7 Localization for the Subtractive Stage
During the subtractive stage of the algorithm, each active robot $u_{n}$ with $n=N+k$ shall moves from lattice point $s_{r}[k]$ to lattice point $t_{r}[k]$. We use $R$ to denote the ribbon that $s_{r}[k]$ belongs to, and recall that we use $R^{[p]}$ to denote the parent ribbon of $R$. In this section, we prove the localizability of the robots of the subtractive stage. Notice that all non-moving robots are localizable at the end of the additive stage by proposition 16 and theorem 18. We are left to show the localizability of every moving robot $u_{n}$ on its path from $s_{r}[k]$ to $t_{r}[k]$.

### v3

@@ B.7 Conventions and Notations
  • robot $u$ belongs to ribbon $R$ if its center $p_{n^{-}}(u)$ is a vertex of $R$, denoted as $u\in R$. The gradient value of robot $u$ is the gradient value of the ribbon $R$;
  • robot $u$ is in front of robot $v$ if $u,v$ belongs to the same ribbon and lattice point $p_{n^{-}}(u)\leftarrow p_{n^{-}}(v)$.
@@
  • a ribbon $R$ is filled if
    $\forall x\in V(R),\exists\text{ non-moving robot }u\in U\text{ such that }x=p_{n^{-}}(u)$
@@ Appendix C Correctness of the Add-Subtract Algorithm
In appendix C, we formalize the add-subtract algorithm and provide a rigorous proof of correctness. The correctness of the problem statement includes two parts: first, we need to show that, for each moving robot, the path between the starting position and stopping position is unobstructed and the moving robot is localizable when needed; and second, we need to show the final coverage of the MRS satisfied the requirement. The second part is can be easily verified by looking at the definition of the assembly sequence and the reassembly sequence. As a result, we shall focus on proving the first part. Suppose we have obtained a proper ribbonization of $S$ and suppose there are $N$ lattice points in $S$. The proof is carried out in three steps.
For the first step, we analyze the additive stage. Firstly, we define the activation sequence $s[k]$ and the assembly sequence $t[k]$ in section C.1. After that, we show the mobility of each moving robot $u_{n}$ by investigating the boundary of the MRS in section C.2. In particular, we prove that the boundary of the MRS maintains a certain smoothness that allows an unobstructed path along the perimeter of the MRS from start position $s[k]$ to target position $t[k]$ for every active robot $u_{k}$ with $k\leq N$. Then, we show the localizability of each moving robot $u_{n}$ in section C.3. In particular, we show that the active robot is able to localize itself at the starting and stopping positions.
For the second step, we investigate the properties of the MRS after the additive stage in section C.4. In particular, we define the "merging point" for every ribbon to denote the lattice point on the perimeter at which a robot moves out of $S$ first visits.
For the third step, we investigate the subtractive stage. We start by formalizing the subtractive stage of the algorithm using the reactivation sequence $s_{r}[k]$ and reassembly sequence $t_{r}[k]$ (section C.5). After that, we investigate the path between lattice point $s_{r}[k]$ and lattice point $t_{r}[k]$, and show that the path for robot $u_{N+k}$ is unobstructed by utilizing the results from the additive stage (section C.6). The localizability of each moving robot is proved in section C.7.
Finally, in section C.8, we conclude that the sequence of tuples defined by the movement sequences satisfies the problem statement.
@@ C.1 Activation Sequence and Assembly Sequence
During the additive stage, in each epoch, one of the idle robots becomes active and moves along the perimeter of the collection of the robots. The moving robot enters the user-defined shape and stops at one of the lattice points inside the shape. In such a way, the MRS shrinks in the idle half-plane $H_{2}$ and expends inside $S$ in the formation half-plane $H_{1}$. We use two sequences of lattice points to represent which idle robot is activated and where it shall stop:
@@ C.2 Path for the Additive Stage
In a similar way, we define the edge-following points and the corresponding edge-following set $\bm{EFS_{n^{+}}}$ for sub-epoch $T=n^{+}$.
@@
We shall prove this by induction. Let $P(n)$ be "at sub-epoch $T=n^{-}$, the set of boundary robots $B_{n^{-}}$ is proper". Let $Q(n)$ be "at sub-epoch $T=n^{+}$, the set of boundary robots $B_{n^{+}}$ is proper".
@@
        Case 3: $v$ is an idle robot. Since any lattice point in $H_{2}$ is at least $2\sqrt{3}d$ from any lattice point in $H_{1}$, any robot occupying a lattice point in $H_{1}$ and an idle robot in $H_{2}$ cannot be adjacent to the same lattice point.
@@
        Case 2 & Case 3: $v$ is a seed robot or an idle robot. By similar arguments as in 1.Case 2 and 1.Case 3, we obtain contradictions.
@@
Since an edge-following path consists of arcs between adjacent edge-following points, we could define a robot’s contribution to the path as follows:
@@ C.3 Localization for the Additive Stage
We shall do induction on the epoch $T$. Let $P(n)$ be "$u_{n}$ is localizable if it neighbors a seed robot or a robot that belongs to a filled ribbon."
@@ C.4 Properties of the MRS At the End of the Additive Stage
    If $x$ has the type $A$ neighborhood, denote the lattice point in the neighborhood with a hop count $k-1$ as $y$. Two lattice points in the neighborhood of $x$ are adjacent to both $x$ and $y$, and we denote the lattice point on the counter-clockwise side of $x$ with respect to $y$ as $a$, and the lattice point on the clockwise side of $x$ as $b$. Since $R$ has a minimum length of two, $a$ is a vertex of $R$ and $a$ is behind $x$. As a result, $b$ is not in $S$, and $b$ is a merging point of $R$. Now, we show that $b$ is the only merging point of $R$.
@@
    Inductive step: suppose $Q(n-1)$ is true, we want to show $P(n)$ is true; and suppose $P(n)$ is true, we want to show that $Q(n)$ is true. Since $s[n]$ is in the idle half-plane $H_{2}$, the non-moving robots in $H_{1}$ and the affiliated path remain unchanged after $u_{n}$ starts to move. As a result, $Q(n-1)\rightarrow P(n)$. We are left to show $P(n)\rightarrow Q(n)$. In other words, we want to show that all the points contributed by robots of $R$ are behind $m$ after $u_{n}$ stops. We shall discuss the following three cases:
@@
First, note that $m$ is on the culminating edge-following path $\widehat{EFP}$ since $m$ is adjacent to a non-moving robot in $S$ at sub-epoch $T=N^{+}$. As a result, $m$ is on $\widehat{AP}$. By proposition 20, the proposition follows.
@@
First, notice that by the definition of merging point, $n$ is contributed by a robot of $R_{i}$. By proposition 21, $m$ is in front of all the points on the affiliated path that are contributed by robots from $R_{i}$. The corollary follows. ∎
@@ C.5 Reactivation Sequence and Reassembly Sequence
During the subtractive stage, extra robots of each layer move outside $S$ while others re-arrange themselves along the layer to fill up the lattice points in $S\setminus D$. The robots moving out of $S$ move along the perimeter of the MRS and join the idle robots in the idle half-plane $H_{2}$. We call such robots "recycled robots". The remaining robots of each layer move within the layer and redistribute to occupy lattice points in $S\setminus D$, and we call such robots "rearranged robots".
@@
The reactivation sequence $s_{r}[k]$ is defined recursively such that each element $s_{r}[k]$ is the smallest lattice point in $S$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not been included in ${s_{r}}_{[1:k-1]}$.
@@
    The rearrangement sequence is a sequence of lattice points in $S\setminus D$, defined recursively such that the $k$th element is the smallest lattice point in $S\setminus D$ (w.r.t the ribbon order) on the largest ribbon (w.r.t the complete tree order) which has not yet been included in the rearrangement sequence.
@@
The re-assembly sequence $t_{r}[k]$ for $1\leq k\leq N$ is generated recursively such that each element $t_{r}[k]$ satisfies:
    if the lattice position $s_{r}[k]$ is among the first $c(R(s_{r}[k]))$ lattice points of the ribbon $R(s_{r}[k])$, $t_{r}[k]$ is the first lattice point in the recycle sequence that has not been include in ${t_{r}}_{[1:k-1]}$;
    if $s_{r}[k]$ is not among the first $c(R(s_{r}[k]))$ lattice points of the ribbon $R(s_{r}[k])$, $t_{r}[k]$ is the first lattice point in the rearrangement sequence that has not been include in ${t_{r}}_{[1:k-1]}$.
Loosely speaking, the robot at the front of the outermost ribbon becomes active first. If the robot is a recycled robot, it moves out of the shape $S$ and stops in $H_{2}$ layer by layer; otherwise, the robot is a rearranged robot and it moves to the front of the ribbon and stops at the smallest unoccupied vertex. This property is captured by the following proposition:
@@
Let sequence $I[k]$ be an index sequence of selects the first vertex of each ribbon in the reactivation sequence $s_{r}$. Notice that $I[k]$ follows the complete tree order in the sense that each element appears earlier in the sequence $I[k]$ is smaller with respect to the complete tree order. Let $Q$ denote the set containing all boundary points of $S$.
@@ C.6.1 Motion of Recycled Robots
For a recycle robot $u_{n}$ with $n=N+k$, suppose $s_{r}[k]$ is a vertex of ribbon $R$ and $R$ has a parent ribbon $R^{[p]}$. The active robot first rotates around the parent ribbon $R^{[p]}$ while keeping a distance $d$ from the nearest robot on $R^{[p]}$ until it reaches the first vertex of $R$. From the first vertex $x$, the robot continues its rotation and reaches the merging point $m(R)$ of the ribbon $R$. Finally, $u_{n}$ moves clockwise along the perimeter of the MRS and stops in $H_{2}$.
@@
Notice that all lattice points in $R$ before $s_{r}[k]$ are unoccupied since all robots that occupied such lattice points have moved out of $S$ and stopped in $H_{2}$ by proposition 23. It is left to show that the curve is unobstructed between each adjacent lattice point on the path. Suppose two adjacent lattice points $x,y$ are vertices of ribbon $R$, the curve between $x$ and $y$ with distance $d$ from robot $v$ (belonging to the parent ribbon of $R$) is shown below.
@@
The extended edge-following set, denoted as $\overline{EFS_{n^{-}}}$, is the set that contains all the extended edge-following points. The same definition applies to the extended edge-following points at $T=n^{+}$ and the corresponding extended edge-following set $\overline{EFS_{n^{+}}}$.
@@
We shall prove this by induction. Let $P(n)$ be "the extended set of boundary robots $\overline{B_{n}^{-}}$ is proper". Let $Q(n)$ be "the extended set of boundary robots $\overline{B_{n}^{+}}$ is proper".
First, notice that the culminating set of boundary robots $\widehat{B}$ is proper. Since the extended set of boundary robots agrees with the culminating set of boundary robots in $H_{1}$, the base case $P(N+1)$ is true.
@@
Recall $p_{3}$ is the lattice point on the right of the origin $O$. Observe that the exterior edge-following path goes through lattice point $p_{3}$. Since virtual robots (if they exist) are in half-plane $H_{1}$, we only need to show that the path segment of the exterior edge-following path from $m(R)$ to $p_{3}$ satisfies the proposition. We denote the path segment from $m(R)$ to $p_{3}$ as $P$. Notice that $P$ is a path segment of the culminating affiliated path $\widehat{AP}$.
By proposition 24, no point on $P$ is contributed by any virtual robot at the time when the first robot of ribbon $R$ starts to move. Let $T=l$ denote the epoch when the first robot of ribbon $R$ starts to move. By the reactivation sequence, all active robots at epoch $l<T<N+k$ are from ribbon $R$. As a result, a point on $P$ is either contributed by a boundary robot or a virtual robot occupying some vertex of $R$. By proposition 21, $m(R)$ is in front of all points of $\widehat{AP}$ that are contributed by robots/virtual robots of $R$. As a result, no point on $P$ (except for the merging point itself) is contributed by a virtual robot. The proposition follows.
@@
For the recycle robot $u_{n}$ with $n=N+k$, its exterior path is well-defined and unobstructed.
@@ C.7 Localization for the Subtractive Stage
During the subtractive stage of the algorithm, each active robot $u_{n}$ with $n=N+k$ shall move from lattice point $s_{r}[k]$ to lattice point $t_{r}[k]$. We use $R$ to denote the ribbon that $s_{r}[k]$ belongs to, and recall that we use $R^{[p]}$ to denote the parent ribbon of $R$. In this section, we prove the localizability of the robots of the subtractive stage. Notice that all non-moving robots are localizable at the end of the additive stage by proposition 16 and lemma 18. We are left to show the localizability of every moving robot $u_{n}$ on its path from $s_{r}[k]$ to $t_{r}[k]$.